# Taylor-Fourier integration

### v1

@@ 4.2.3 High precision numerical propagation of a geostationary orbit
We now consider the initial state of a geostationary satellite [12, pg. 116]. The initial position coordinates (in kilometers) and velocity vectors (in kilometers per second) are
@@
We have computed several $(M,d)$-Taylor-Fourier approximations to the solution of the VOP formulation (44), and found that $M=8$ and $d=8$ is a good choice in that particular initial value problem. For comparison purposes, we have solved (44) numerically (with absolute and relative tolerance set to $10^{-13}$) for 400 revolutions with an explicit RK method of order 9 due to Verner [18], efficiently implemented with adaptive time-stepping as algorithm Vern9 in DifferentialEquations.jl. We have estimated the errors of both numerical approximations by comparing them with a reference solution computed (in high precision floating point arithmetic) with Vern9 with absolute and relative tolerance set to $10^{-20}$.
Figure 5: Relative errors in position (left) and absolute errors in physical time (right) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the numerical integration with Vern9 with tol=$10^{-13}$ (top) and the $(8,8)$-Taylor-Fourier approximation (bottom).
In Figure 5 we display the relative errors in position ${\bf q}$ (left subplot) and the absolute errors in physical time $t$ (right subplot) versus the fictitious time $\tau$ scaled by the period $P=2\pi/\omega$ of the satellite orbits. The upper (resp. lower) plots correspond to errors of Vern9 (resp. Taylor-Fourier). Observe that the absolute errors in physical time are smaller for the Taylor-Fourier approximation. Moreover, while similar relative errors in position are obtained for both approximations at the end of the integration interval, computing the Taylor-Fourier approximation required considerably less CPU time than Vern9 (by a factor of 20 on our laptop).
We emphasize that while Vern9 exhibits the expected pattern of increasing position errors due to the propagation of local truncation and round-off errors, the Taylor-Fourier approximation maintains positional errors below a threshold of approximately $3\times 10^{-15}$ (corresponding to the errors due to truncation of Fourier modes) for approximately 380 revolutions. Using $M=16$ (not shown here) instead of $M=8$, that threshold lowers to round-off levels. Observe that after $380$ revolutions, these errors start to grow. We have checked that error growth is approximately proportional to $t^{d+1}$. The validity of the Taylor-Fourier approximation can be extended to longer time intervals by increasing the degree $d$. For instance, with $d=9$, the errors in position remain below the above-mentioned threshold for up to 500 revolutions.
@@ 4.2.4 High precision numerical propagation of a highly eccentric orbit
We now make similar computations with the initial state of a highly eccentric orbit (with eccentricity $e=0.7679436$) obtained from [12, pg. 51]. In particular,
@@
Figure 6: Relative errors in position (left) and absolute errors in physical time (right) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the numerical integration with Vern9 with tol=$10^{-13}$ (top) and the $(128,14)$-Taylor-Fourier approximation (bottom).
@@
Figure 6 is similar to Figure 5, the only differences being that we display the errors for up to 40 revolutions (instead of 400) and that $(M,d)=(128,14)$ (instead of $(M,d)=(8,8)$). Vern9 is still applied with absolute and relative tolerances set to $10^{-13}$.
As in the previous example, the absolute errors in physical time are smaller for the Taylor-Fourier approximation. In this second example, the relative errors in position are also smaller for the Taylor-Fourier approximation, while the computation of the Taylor-Fourier approximation required considerably less CPU time than Vern9 (by a factor of $2.6$ on our laptop).
@@ 5 Conclusions
We have proposed approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions that after an appropriate change of variables can be written as a non-autonomous system with $(2\pi/\omega)$-periodic dependence on $t$. These approximate solutions, that can be written in closed form combining truncated Fourier and Taylor series, are valid for a relatively large time interval and, once computed, can be efficiently evaluated at arbitrary times in that interval. The resulting approximate solutions are uniformly accurate in the basic frequency $\omega$, which makes them more efficient than standard numerical integrators when $\omega$ is large enough. In addition, we have provided a detailed procedure to compute efficiently such approximate solutions making use of power series arithmetic techniques and the FFT algorithm.
We have also reported numerical results to show the performance of the Taylor-Fourier approximations when compared with state-of-the-art methods. When applied to a perturbed Kepler problem, the proposed approximations become more efficient than the numerical integration of the transformed system using the adaptive algorithm Vern9, even for highly eccentric orbits. The application of the Taylor-Fourier integration procedure to the semi-discretized cubic nonlinear Schrödinger equation is not so advantageous from the computational point of view when compared with the Lawson integrator based on the variable-step implementation of DP5. This is mainly due to the high dimension of the semi-discretized problem. However, the uniform accuracy of Taylor-Fourier approximations with respect to the basic frequency $\omega$ implies that they are more efficient than standard numerical methods (which typically experience accuracy deterioration as the frequency $\omega$ increases) for high enough values of $\omega$.
A detailed error analysis of Taylor-Fourier approximations, and their connection with other related topics such as stroboscopic averaging and the solution of the transport equation will be the subject of future work.

### v2

@@ 6.2.3 High precision numerical propagation of a geostationary orbit
We now consider the initial state of a geostationary satellite [14, pg. 116]. The initial position coordinates (in kilometers) and velocity vectors (in kilometers per second) are
@@
We have computed several $(M,d)$-Taylor-Fourier approximations to the solution of the VOP formulation (67) over 400 revolutions, and found that $M=8$ and $d=8$ is a good choice in that particular initial value problem. We have estimated the errors of the Taylor-Fourier approximations by comparing them with a reference solution computed (in high precision floating point arithmetic) with Vern9, an explicit RK method of order 9 due to Verner [20], efficiently implemented with adaptive time-stepping in DifferentialEquations.jl. The absolute and relative tolerances have been set equal to $10^{-20}$.
Figure 4: Relative errors in position (top) and absolute errors in physical time (bottom) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the $(8,8)$-Taylor-Fourier approximation.
In Figure 4 we display the relative errors in position ${\bf q}$ (top subplot) and the absolute errors in physical time $t$ (bottom subplot) versus the fictitious time $\tau$ scaled by the period $P=2\pi/\omega$ of the satellite orbits for the $(8,8)$-Taylor-Fourier approximation.
We observe that the Taylor-Fourier approximation maintains positional errors below a threshold of approximately $3\times 10^{-15}$ (corresponding to the errors due to truncation of Fourier modes) for approximately 380 revolutions. Using $M=16$ (not shown here) instead of $M=8$, that threshold lowers to round-off levels. Observe that after $380$ revolutions, these errors start to grow. We have checked that error growth is approximately proportional to $t^{d+1}$. The validity of the Taylor-Fourier approximation can be extended to longer time intervals by increasing the degree $d$. For instance, with $d=9$, the errors in position remain below the above-mentioned threshold for up to 500 revolutions.
Figure 5: Energy error of the averaging approximated solution for $(M,d)=(8,8)$ (top) and the position error of the averaged approximation with respect the $(8,8)$-Taylor-Fourier approximation (bottom) in 600 revolutions.
We next check the accuracy of averaged approximation on the left-hand side of (41). We show in Figure 5 the energy error of the averaged approximated solution (top subplot) and the relative error of the position with respect to the Taylor-Fourier approximation (bottom). It can be observed that the behavior of the averaged approximation is practically identical to the behavior of the Taylor-Fourier approximation: it maintains the energy error at rounding error levels for approximately 380 revolutions, just like the Taylor-Fourier approximation, and on the other hand, the difference between the averaged approximation and the Taylor-Fourier approximation remains below the rounding error level for more than 600 revolutions.
@@ 6.2.4 High precision numerical propagation of a highly eccentric orbit
We now make similar computations with the initial state of a highly eccentric orbit (with eccentricity $e=0.7679436$) obtained from [14, pg. 51]. In particular,
@@
Figure 6: Relative errors in position (top) and absolute errors in physical time (bottom) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the $(128,14)$-Taylor-Fourier approximation.
Figure 6 is similar to Figure 4, the only differences being that we display the errors for up to 40 revolutions (instead of 400) and that $(M,d)=(128,14)$ (instead of $(M,d)=(8,8)$).
@@ 7 Conclusions
We have proposed approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions that after an appropriate change of variables can be written as a non-autonomous system with $(2\pi/\omega)$-periodic dependence on $t$. These approximate solutions, that can be written in closed form combining truncated Fourier and Taylor series, are valid over a time interval whose length is independent of $\omega$. Once computed, they can be efficiently evaluated at arbitrary times within that interval. The resulting approximate solutions are uniformly accurate [6] with respect to the basic frequency $\omega$.
We have provided a detailed procedure to compute efficiently such approximate solutions making use of power series arithmetic techniques and the FFT algorithm. Additionally, we have demonstrated how Taylor-Fourier approximations can be used to evaluate various maps associated with high-order averaging of highly oscillatory systems. Furthermore, we have extended the Taylor-Fourier approximation methodology to highly oscillatory ODEs with multiple non-resonant frequencies.
We have also reported numerical experiments to illustrate the application of the Taylor-Fourier approximations. Their uniform accuracy with respect to the basic frequency $\omega$ suggests that they could be more efficient than standard numerical methods (which typically suffer from accuracy deterioration as $\omega$ increases) for sufficiently high values of $\omega$. When applied to a perturbed Kepler problem (the so-called $J_{2}$-problem in orbit propagation), the proposed approximations efficiently provide very accurate solutions, even for highly eccentric orbits. The Taylor-Fourier approximations to the solution of the semi-discretized cubic nonlinear Schrödinger equation reproduce the well-known Talbot phenomenon, though at the cost of increased computational effort due to the high dimensionality of the semi-discretized problem.